# A Family of Convex Models to Achieve Fairness through Dispersion Control

### v1

@@ Abstract
Controlling the dispersion of a subset of decision variables in an optimization problem is crucial for enforcing fairness or load-balancing across a wide range of applications. Building on the well-known equivalence of finite-dimensional norms, the note develops a family of parameterized convex models that regulate the dispersion of a vector of decision-variable values through its coefficient of variation. Each model contains a single parameter that takes a value in the interval $[0,1]$. When the parameter is set to zero, the model imposes only a trivial constraint on the optimization problem; when set to one, it enforces equality of all the decision variables. As the parameter varies, the coefficient of variation is provably bounded above by a monotonic function of that parameter. The note also presents theoretical results that relate the space of feasible solutions to all the models.
1 Introduction
@@
Fairness has deep roots in game theory, where foundational bargaining solutions were rigorously characterized via axiomatic frameworks. Nash’s seminal work formalized the bargaining problem through a set of axioms–Pareto efficiency, symmetry, invariance to affine transformations, and independence of irrelevant alternatives–which uniquely identify the Nash bargaining solution nash1950bargaining . Subsequent refinements, such as the Kalai-Smorodinsky solution, altered these axioms to prioritize proportionality among agents kalai1977proportional . In economics and communication theory xinying2023guide ; shakkottai2008network , these game-theoretic fairness axioms guide the design of utility functions so that decentralized, self-interested agents naturally converge to equitable outcomes. By carefully embedding fairness criteria into each agent’s payoff, one ensures that the collective solution respects core notions of equity and parity, even when each participant acts independently.
@@
However, in many application domains, problems are centrally optimized: a single decision maker minimizes a global cost or maximizes overall efficiency while imposing load-balancing requirements to distribute burdens–battery discharge, travel distance, computational load ghomi2017load ; janiar2021deep ; bhadoriya2024equitable –uniformly across agents. Here, “fairness” is more accurately viewed as dispersion control, wherein the vector of decision-variable values must exhibit limited spread under a centralized objective. In contemporary optimization, efforts have been made to extend beyond achieving minimal cost or maximal utility by explicitly regulating the dispersion of selected decision variables. By bounding or penalizing the spread through metrics such as variance, range, coefficient of variation, or more sophisticated equity indices xinying2023guide , one can embed fairness or load-balancing constraints directly into existing optimization models. However, many dispersion metrics do not admit convex reformulations; directly bounding them thus renders the optimization problem non‑convex and computationally challenging xinying2023guide . Nevertheless, the dispersion‑control paradigm is ubiquitous: ensuring equitable service levels in cloud-computing clusters ghomi2017load , allocating shared spectrum in wireless networks matyjas2016spectrum , operating power grids during extreme events sundar2025parametric , among others. Controlling dispersion thus provides a unifying formalism that couples classical efficiency criteria with societally critical notions of parity, equity, and fairness.
1.1 Literature Review
In the literature, two principal modelling paradigms have emerged for incorporating fairness into decision-making problems that traditionally focus on maximizing efficiency.
[…]
Fairness levels are typically assessed a posteriori using indices such as the Gini coefficient gini1936measure , Atkinson index atkinson1970measurement , Hoover index long1997hoover , and Jain’s index jain1984quantitative etc. A substantial body of work has sought to enforce fairness by directly minimizing or maximizing such indices, or by reducing dispersion measures such as the coefficient of variation or the range of decision variables xinying2023guide . These approaches, however, generally suffer from one of two drawbacks: (i) the resulting model is non-convex, or (ii) the model produces a single fair solution that compromises efficiency–an outcome often unacceptable in applications where cost considerations are critical. Despite these issues, this paradigm remains widely used, given the limited availability of alternative frameworks for fairness in centralized optimization.
The second, more recent paradigm–which motivates the present study–retains an efficiency-centric objective while constraining dispersion. Examples include the $\varepsilon$-fairness second-order cone approach sundar2025parametric or the imposition of explicit upper bounds on fairness indices such as the Gini coefficient bektacs2020using , Jain’s index, coefficient of variation or range xinying2023guide . Here, fairness is modelled by enforcing a lower (or upper) bound on the minimum (or maximum) acceptable fairness level while preserving the efficiency objective. This approach does not always yield a convex optimization problem, as noted in xinying2023guide , and even when convexity is preserved, computing a suitable lower bound for fairness often requires trial-and-error. Two notable exceptions are the $\varepsilon$-fairness second-order cone method sundar2025parametric , which achieves this using a single second-order cone constraint, and the direct bounding of the Gini coefficient bhadoriya2024equitable , which reformulates the problem into a set of linear inequalities at the cost of introducing $O(n^{2})$ auxiliary variables. For a comprehensive review of both paradigms, the reader is referred to the survey in xinying2023guide . This note examines $\varepsilon$-fairness and extends it to a family of convex models. The appeal of $\varepsilon$-fairness lies in its desirable properties: it enforces fairness through a single parametric convex constraint with $\varepsilon\in[0,1]$ as the parameter, and the parameter $\varepsilon$ is monotonic in the level of fairness. Moreover, it naturally facilitates the study of the Pareto front of optimal solutions balancing efficiency and fairness in a bi-objective setting bhadoriya2024equitable .
We now present the main contributions of this note in the context of this literature.
1.2 Contributions
@@
The contributions of the note are as follows: we extend the recently introduced model of $\varepsilon$-fairness sundar2025parametric –that controls the dispersion of decision variables via its coefficient of variation everitt2010cambridge –to a countably infinite family of convex models. Each model is derived from the first principles through the classical equivalence of finite-dimensional norms halmos2017finite , namely between the $\ell_{p}$-norm and the $\ell_{1}$-norm for any $p\geqslant 2$. All models are governed by a single parameter, $\varepsilon\in[0,1]$: when $\varepsilon=0$, no constraint on the dispersion is imposed, and when $\varepsilon=1$, the constraint forces the dispersion to be zero. In particular, for $p=2$, the constraint reduces to a second-order cone sundar2025parametric ; for $p=\infty$, it reduces to a set of linear inequalities; and for $p\in(2,\infty)$, it reduces to the $\ell_{p}$-norm cone. In addition, we present theoretical results that characterize the relationships among these convex models for all values of $\varepsilon$, including their relative strength, inclusion properties, and points of equivalence.
2 Model Development
@@
We begin by introducing notation and recalling a classical result on the equivalence of norms in finite-dimensional spaces halmos2017finite . Let $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ denote an $n$-dimensional non-negative vector, whose dispersion we wish to regulate. For any integer $p\geqslant 1$, define its $\ell^{p}$‐norm by
@@
Here we omit the absolute value in (1) since $x_{i}\geqslant 0$ for all $i$. Also define
@@
Then, for any pair of integers $p_{1}$ and $p_{2}$, such that $0<p_{1}<p_{2}$, the norm‐equivalence result halmos2017finite states
$\mathcal{NE}(p_{1},p_{2}):\;\;\|\bm{x}\|_{p_{2}}\;\leqslant\;\|\bm{x}\|_{p_{1}}\;\leqslant\;\left(\frac{D_{p_{2}}+1}{D_{p_{1}}+1}\right)\,\|\bm{x}\|_{p_{2}}.$ (3)
@@
  • The lower bound $\|\bm{x}\|_{p}=\|\bm{x}\|_{1}$ is attained if and only if all but one component of $\bm{x}$ are zero (the most unfair distribution of $\bm{x}$).
  • The upper bound $\|\bm{x}\|_{1}=(D_{p}+1)\,\|\bm{x}\|_{p}$ is attained if and only if all components of $\bm{x}$ are equal (the most fair distribution of $\bm{x}$).
This intuition naturally leads to the following definition, extending the concept of $\varepsilon$-fairness in sundar2025parametric .
Definition 1
@@
Furthermore, we say that $\bm{x}$ is at least $(\varepsilon,p)$‐fair if
@@
In the above definition, the factor $(1+\varepsilon D_{p})$ can be interpreted as the convex combination of $1$ and $(D_{p}+1)$, with $\varepsilon$ as the non-negative multiplier. Furthermore, the constraint in (5) is convex for every $p\geqslant 2$. In particular, for $p=2$, (5) reduces to a second‐order cone sundar2025parametric ; for $2<p<\infty$, it defines an $\ell_{p}$‐norm cone, and for $p=\infty$, it reduces to a collection of linear inequalities as follows:
$\|\bm{x}\|_{1}\;\geqslant\;(1+\varepsilon D_{\infty})\,\|\bm{x}\|_{\infty}\iff\sum_{i=1}^{n}x_{i}\;\geqslant\;\left(1+n\varepsilon-\varepsilon\right)\cdot x_{i}\quad\forall\,i.$ (7)
@@
The following normalization argument to simplifies notation for the main theoretical results in the subsequent section.
Remark 1
Normalization – Since all finite-dimensional norms are positively homogeneous of degree 1, the inequality in (6), i.e.,
@@
is invariant under scaling of $\bm{x}$. Thus, it is convenient to normalize any non-zero vector [^1] The zero vector, $\bm{0}$, cannot be normalized and is trivially $(\varepsilon,p)$-fair for any $p\geqslant 2$ and $\varepsilon\in[0,1]$. Hence, throughout the rest of the note, we assume that $\bm{x}$ is always a non-zero vector. by its $\ell_{1}$-norm. To that end, define
@@
3 Theoretical Properties
We first establish the relationship between the $(\varepsilon,p)$-fairness of Definition 1 and the coefficient of variation of the vector $\bm{x}$. The coefficient of variation of $\bm{x}$ is a statistical measure of the dispersion of $\bm{x}$. The result in the subsequent section 3.1 provides the intuition for how varying $\varepsilon$ in (5) for any fixed integer $p\geqslant 2$ allows the control of dispersion of $\bm{x}$.
3.1 Bound of Coefficient of Variation
@@
$\mathrm{CV}(\bm{x})\triangleq\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}}}{\mu},\quad\text{ where }\quad\mu=\frac{1}{n},$
@@
On the other hand, since $n\mu=1$ and observing that
@@
Finally, since $D_{p}>0$, we have
@@
Hence, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ and $B_{p}(0)=(D_{p}+1)^{2}-1$, $B_{p}(1)=0$, as claimed.
The strict monotonicity of $B_{p}(\varepsilon)$ in $\varepsilon$ implies that, as $\varepsilon$ increases, the upper bound on $\mathrm{CV}(\bm{x})$ decreases, thereby decreasing the dispersion of $\bm{x}$. This relationship was previously established for the case $p=2$ in sundar2025parametric .
In the forthcoming sections, we present the main theoretical results that concern the equivalence or ordered relation between sets $\mathcal{Y}(p,\varepsilon)$ and correspondingly of $\mathcal{X}(p,\varepsilon)$ for varying values of $p$.
3.2 Equivalence Condition
@@
Proof
The proof follows from the observation in Section 2 that the term $(1+\varepsilon D_{p})$ is the convex combination of the two scalars $1$ and $(D_{p}+1)$ with $\varepsilon$ as the factor.
Theorem 3.1
@@
When $\varepsilon=1$, Proposition 3 indicates that $\mathcal{Y}(1,p)=\left\{\tfrac{1}{n}\bm{e}\right\}$ for every $p\geqslant 2$.
Theorem 3.1 shows that at the two extreme fairness levels, the feasible set of vectors $\bm{x}$ is the same regardless of the choice of $p$.
3.3 Strict Inclusion Property
@@
Proof
See Appendix A.
Theorem 3.2
@@
$(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1\iff\varepsilon\leqslant\varepsilon_{p}(\bm{x})$
@@
The excluded cases correspond to the boundary situation $\varepsilon\in\{0,1\}$, already treated in Proposition 3 and Theorem 3.1. For any such $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, the proof is complete if we show that the map

### v2

@@ Abstract
Controlling the dispersion of a subset of decision variables in an optimization problem is crucial for enforcing fairness or load-balancing across a wide range of applications, from robotics to communication networks, supply-chain management, as well as public policy design. Building on the well-known equivalence of finite-dimensional norms, we develop a family of parameterized convex models that regulate the dispersion of a vector of decision-variable values through its coefficient of variation. Each model contains a single parameter defined on the interval $[0,1]$. When the parameter is set to 0, the model imposes only a trivial constraint on the optimization problem; when set to 1, it forces all decision variable values to become equal. As the parameter varies, the coefficient of variation is provably bounded above by a monotonic function of that parameter. We present theoretical results that relate the feasibility space of all the models and compare the models’ computational time and solution quality on a multi-agent assignment problem. Finally, we discuss the implications of these models for optimization problems in other domains.
I Introduction
@@
Fairness has deep roots in game theory, where foundational bargaining solutions were rigorously characterized via axiomatic frameworks. Nash’s seminal work formalized the bargaining problem through a set of axioms–Pareto efficiency, symmetry, invariance to affine transformations, and independence of irrelevant alternatives–which uniquely identify the Nash bargaining solution [1]. Subsequent refinements, such as the Kalai-Smorodinsky solution, altered these axioms to prioritize proportionality among agents [2]. In economics and communication theory [3, 4], these game-theoretic fairness axioms guide the design of utility functions so that decentralized, self-interested agents naturally converge to equitable outcomes. By carefully embedding fairness criteria into each agent’s payoff, one ensures that the collective solution respects core notions of equity and parity, even when each participant acts independently.
However, in many application domains like control, robotics, and logistics, problems are centrally optimized: a single decision maker minimizes a global cost or maximizes overall efficiency while imposing load-balancing requirements to distribute burdens–battery discharge, travel distance, computational load [5, 6, 7]–uniformly across agents. Here, “fairness” is more accurately viewed as dispersion control, wherein the vector of decision-variable values must exhibit limited spread under a centralized objective.
@@
In contemporary optimization, efforts have been made to extend beyond achieving minimal cost or maximal utility by explicitly regulating the dispersion of selected decision variables. By bounding or penalizing the spread through metrics such as variance, range, coefficient of variation, or more sophisticated equity indices [3], one can embed fairness or load-balancing constraints directly into existing optimization models. However, many dispersion metrics do not admit convex reformulations; directly bounding them thus renders the optimization problem non‑convex and computationally challenging [3]. Nevertheless, the dispersion‑control paradigm is ubiquitous: ensuring equitable service levels in cloud-computing clusters [5, 8], allocating shared spectrum in wireless networks [6, 9], managing electric-vehicle fleet loads [7], operating power grids during extreme events [10], and coordinating distributed energy resources [11], among others.
Controlling dispersion thus provides a unifying formalism that couples classical efficiency criteria with societally critical notions of parity, equity, and fairness. In the literature, two principal modeling paradigms have emerged. The first replaces the efficiency objective with an explicitly fairness-oriented objective–e.g., Rawlsian max–min fairness [12], $\alpha$-fairness [13], proportional fairness [4], or $\ell_{p}$-norm dispersion objectives [14]–so that parity is pursued as the primary goal. The second, which motivates the present study, retains an efficiency-centric objective while constraining dispersion, for example via the $\varepsilon$-fairness second-order cone [10] or explicit upper bounds on canonical indices such as the Gini coefficient [15, 14] and the Jain-index [16, 3]. The exhaustive survey in [3] reviews both approaches, and recent applications of the “efficiency-subject-to-fairness” paradigm span vehicle routing [7], power-grid operations and control [10, 11], and inventory control [17].
@@
The contributions of the article are as follows: we extend the recently introduced model of $\varepsilon$-fairness [10]–that controls the dispersion of decision variables via its coefficient of variation [18]–to a countably infinite family of convex models. Each model is derived from the first principles through the classical equivalence of finite-dimensional norms [19], namely between the $p$-norm and the $1$-norm for any $p\geqslant 2$. All models are governed by a single parameter, $\varepsilon\in[0,1]$: when $\varepsilon=0$, no constraint on the dispersion is imposed, and when $\varepsilon=1$, the constraint forces the dispersion to be zero. In particular, for $p=2$, the constraint reduces to a second-order cone [10]; for $p=\infty$, it reduces to a set of linear inequalities; and for $p\in(2,\infty)$, it yields the familiar family of $\ell_{p}$-norm cones. The article also provides a comprehensive set of theoretical results that characterize the relationships among these convex models for all values of $\varepsilon$, detailing their relative strength, inclusion properties, and points of equivalence. Finally, we compare and contrast the different models in terms of computation time and its effectiveness in controlling dispersion on a multi-agent task assignment problem–an important problem that arises in many civilian and military applications like target tracking [20], intelligent transportation [21] and logistics [22], and search and rescue [23]. In particular, we formulate a variant of the assignment problem that aims to control the dispersion of the total time consumed to complete the tasks across different agents.
The rest of the article is organized as follows: Section II develops the fairness models from first principles, followed by theoretical properties of the models in Section III. Section IV formulates the fairness-constrained multi-agent assignment problem to demonstrate the use of the fairness models. Computational results that corroborate the effectiveness of the fairness models are presented in Section V followed by concluding remarks in Section VI.
II Model Development
@@
We begin by introducing notation and recalling a classical result on the equivalence of norms in finite-dimensional spaces [19]. Let $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ denote an $n$-dimensional non-negative vector, whose dispersion we wish to regulate. For any integer $p\geqslant 1$, define its $\ell^{p}$‐norm by
@@
Here we omit the absolute value in (1) since $x_{i}\geqslant 0$ for all $i$. Define also
@@
Then, for any pair of integers $p_{1}$ and $p_{2}$, such that $0<p_{1}<p_{2}$, the norm‐equivalence result [19] states
$\mathcal{NE}(p_{1},p_{2}):\;\;\|\bm{x}\|_{p_{2}}\;\leqslant\;\|\bm{x}\|_{p_{1}}\;\leqslant\;\frac{D_{p_{2}}+1}{D_{p_{1}}+1}\,\|\bm{x}\|_{p_{2}}.$ (3)
@@
  • The lower bound $\|\bm{x}\|_{p}=\|\bm{x}\|_{1}$ is attained if and only if all but one component of $\bm{x}$ are zero (the most unfair distribution).
  • The upper bound $\|\bm{x}\|_{1}=(D_{p}+1)\,\|\bm{x}\|_{p}$ is attained if and only if all components of $\bm{x}$ are equal (the most fair distribution).
This intuition naturally leads to the following definition, extending the concept of $\varepsilon$-fairness in [10].
Definition 1.
@@
Thus, $\varepsilon=0$ corresponds to the most unfair distribution, and $\varepsilon=1$ to the perfectly fair distribution. Equivalently, we say that $\bm{x}$ is at least $(\varepsilon,p)$‐fair if
@@
In the above definition, the factor $(1+\varepsilon D_{p})$ can be interpreted as the convex combination of the endpoints $1$ and $(D_{p}+1)$, with $\varepsilon$ as the non-negative multiplier. Furthermore, the constraint in (5) is convex for every $p\geqslant 2$. In particular:
  • For $p=2$, (5) reduces to a second‐order cone [10].
  • For $2<p<\infty$, it defines an $\ell_{p}$‐norm cone.
  • For $p=\infty$, it reduces to a collection of linear inequalities as follows:
    $\|\bm{x}\|_{1}\;\geqslant\;(1+\varepsilon D_{\infty})\,\|\bm{x}\|_{\infty}$
    $\;\Longleftrightarrow\;$ $\sum_{i=1}^{n}x_{i}\;\geqslant\;\left(1+n\varepsilon-\varepsilon\right)x_{i}\quad\forall\,i=1,\dots,n.$ (7)
@@
We now introduce a normalization argument to simplify notation, before stating and proving our main results.
Remark 1.
Normalization – Since all finite-dimensional norms are positively homogeneous of degree 1 [19], the inequality in (6), i.e.,
@@
is invariant under scaling of $\bm{x}$. Thus, it is convenient to normalize any non-zero vector [^1] The zero vector, $\bm{0}$, cannot be normalized and is trivially $(\varepsilon,p)$-fair for any $p\geqslant 2$ and $\varepsilon\in[0,1]$. Hence, throughout the rest of the article, we assume that $\bm{x}$ is always a non-zero vector. by its $\ell_{1}$-norm. To that end, define
@@
III Theoretical Properties
We first establish the relationship between the $(\varepsilon,p)$-fairness of Definition 1 and the coefficient of variation of the vector $\bm{x}$. The coefficient of variation of $\bm{x}$ is a statistical measure of the dispersion of $\bm{x}$. This result provides the intuition for how varying $\varepsilon$ in (5) for any fixed integer $p\geqslant 2$ allows us to control of dispersion of $\bm{x}$.
III-A Bound of Coefficient of Variation
@@
$\mathrm{CV}(\bm{x})=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}}}{\mu},\quad\mu=\frac{1}{n},$
@@
On the other hand, since $n\mu=1$ and observing
@@
Finally, since $D_{p}>0$, one computes
@@
Hence, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ and $B_{p}(0)=(D_{p}+1)^{2}-1$, $B_{p}(1)=0$, as claimed. ∎
The strict monotonicity of $B_{p}(\varepsilon)$ in $\varepsilon$ implies that, as $\varepsilon$ increases, the upper bound on $\mathrm{CV}(\bm{x})$ decreases, thereby decreasing the dispersion of $\bm{x}$. This relationship was previously established for the case $p=2$ in [10].
In the forthcoming sections, we present the main theoretical results of this article on the equivalence or ordered relation between sets $\mathcal{Y}(p,\varepsilon)$ and correspondingly of $\mathcal{X}(p,\varepsilon)$ for varying values of $p$.
III-B Equivalence Condition
@@
Proof.
The proof follows from the observation in Section II that the term $(1+\varepsilon D_{p})$ is the convex combination of the two scalar values $1$ and $(D_{p}+1)$ with $\varepsilon$ as the factor. ∎
Theorem 1.
@@
When $\varepsilon=1$, Proposition 3 indicates that $\mathcal{X}(1,p)=\tfrac{1}{n}\bm{e}$ for every $p\geqslant 2$. ∎
Theorem 1 shows that at the two extreme fairness levels, the feasible set of vectors $\bm{x}$ is the same regardless of the choice of $p$.
III-C Strict Inclusion Property
@@
Proof.
See Appendix -A. ∎
Theorem 2.
@@
$(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1,\iff\varepsilon\leqslant\varepsilon_{p}(\bm{x}),$
@@
The excluded cases correspond to the boundary situation $\varepsilon\in\{0,1\}$, already treated in Proposition 3 and Theorem 1. For any such $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, the proof is complete if we show that the map